OpenStudy (anonymous):

At a certain time during the day a tall tree casts a shadow of 32 feet. A shorter tree 12 feet tall casts a shadow of 16 feet. What is the height of the taller tree? 28 feet 24 feet 20 feet 16 feet

OpenStudy (btaylor):

Since the sun is in the same spot, it creates similar triangles:  What is the ratio you will use?

OpenStudy (anonymous):

32=12:16 ?

OpenStudy (btaylor):

almost. h:32=12:16.

OpenStudy (anonymous):

oh.. haha I ment that. so then I multiply 32 by 12 then divide by 16?

OpenStudy (anonymous):

H=24

OpenStudy (nurali):

The angle of the sun is the same for both, so the ratio of height to shadow-length is the same for both trees. Thus, if h is the height of the taller tree, h/32 = 12/16 h = 32 (12/16) = 24 feet

OpenStudy (btaylor):

yep.